Your sister just deposited $11,000 into an investment account. She believes that she will earn an annual return of 9.9 percent for the next 6 years. You believe that you will only be able to earn an annual return of 9 percent over the same period. How much more must you deposit today in order to have the same amount as your sister in 6 years?
Answer #1

First, we need to find the future value of the sister's investment:

FV = PV(1 + r)n = $11,000 x 1.0996 = $19,381.12

Now, we can find the deposit needed today to match the FV of the sister's investment:

PV = FV/(1 + r)n = $19,381.12 / 1.096 = $11,556.33

Additional Deposit Needed = $11,556.33 - $11,000 = $556.33